plane b is flying 55 mph faster than plane a. Find the time it takes for plane b to travel 2220 miles if it takes plane a the same amount of time to travel 2000 miles.

Found 2 solutions by TimothyLamb, richwmiller:
Answer by TimothyLamb(4379)

b = a + 55
s = d / t
t = d / s
2000/a = 2220/b
---
2000/a = 2220/(a + 55)
(1/2000)a = (1/2220)(a + 55)
(1/2000)a = (1/2220)a + (1/2220)55
(1/2000)a - (1/2220)a = (1/2220)55
(2220/(2220*2000))a - (2000/(2220*2000))a = (1/2220)55
(220/(2220*2000))a = (1/2220)55
a = 55(2220*2000)/(220*2220)
a = 55(2000)/(220)
a = 55(2000)/(220) = 454.545
---
Answer:
a = 454.545 mph
b = 504.545 mph
time = 4.4 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by richwmiller(17219)

(a+55)*t=2220,
a*t=2000
(a+55)*t=2220,
a = 500 mph, t = 4 hours
b=555 mph